OpenStudy (anonymous):

justin has twice as much as Scott and Scott has 11 more than bobby, together they have $49. How much money does each have?

OpenStudy (lucaz):

let J = Justin's money, S = Scott's money and B = Bobby's money together they have $49, so J+S+B=49 Scott has $11 more than Bobby, S = B+11 Justin has twice as much as Scott, J = 2*S = 2(B+11) substituting.. 2(B+11) + B+11 + B = 49

OpenStudy (lucaz):

solve for B, substitute it back to find S and J